What is the slope of a line perpendicular to 3x+2y=7?
Answer by mananth(16946)

3x+2y=7
..
add -3x
3x-3x+2y=7-3x
2y= -3x+7
y -3/2 x +7/2
..
the slope of this line = -3/2
slope of perpendicular line will be
negative reciprocal
= 2/3